OpenStudy (anonymous):

A large emerald with a mass of 378.24 grams was recently discovered in a mine. If the density of the emerald is 3.91 grams over centimeters squared, what is the volume?

OpenStudy (australopithecus):

set up a ratio 378.24g/x cm^3 = 3.91g/1cm^3 now solve for x

OpenStudy (anonymous):

I don't know how...

OpenStudy (australopithecus):

x = the volume 378.24g takes up

OpenStudy (australopithecus):

are you unfamiliar with basic algebra? If so I can tell you the rules you need to apply?

OpenStudy (anonymous):

I was never good at algebra.

OpenStudy (australopithecus):

to solve the problem, I'm just making the assumption that the density is equal to 378.24g/x cm^3 of the material, this makes sense based on the simple rule: \[\frac{x}{y} = \frac{a*x}{a*y}\] where, \[\frac{a}{a} = 1\] where, a = any number other then 0

OpenStudy (australopithecus):

so for example. \[\frac{2}{4} = \frac{2*5}{4*5} = 1*\frac{2}{4}\] because 5/5 = 1

OpenStudy (australopithecus):

do you understand how I came up with this now?

OpenStudy (australopithecus):

in regards to the formula I posted above?

OpenStudy (anonymous):

Yes

OpenStudy (australopithecus):

Ok awesome, \[\frac{y}{x} = \frac{a}{b}\] \[\frac{y}{x}*\frac{x}{1} = \frac{a}{b}*\frac{x}{1}\] \[\frac{y*x}{x*1} = \frac{a*x}{b*1}\] \[\frac{y*x}{x} = \frac{a*x}{b}\] \[y*1 = \frac{a*x}{b}\] \[y = \frac{ax}{b}\] this property works with division also, the denominator is 1 because you can express any number with a denominator 1 also I wanted to demonstrate basic fraction multiplication rule. Do the same with b, then multiply both sides of the equation by (1/a) and you will have solved for x

OpenStudy (australopithecus):

this is essentially how you will solve this problem remember a, b, y are all just numbers just like your problem

OpenStudy (anonymous):

I'm confused again....I still don't know how to solve the problem.

OpenStudy (australopithecus):

\[\frac{378.24g}{x cm^3} = \frac{3.91g}{1cm^3}\]

OpenStudy (australopithecus):

does this make it easier?

OpenStudy (australopithecus):

you need solve for x to find the volume of 378.24grams of this material. First step you want to bring x up out of the denominator

OpenStudy (australopithecus):

so just multiply both sides by (x/1)

OpenStudy (anonymous):

Yes that makes it easier, but I'm still confused on how to find out what x is. Would I divide 378.24g by 3.91 g to get 96.7 And that would be x?

OpenStudy (australopithecus):

just write it out like I did

OpenStudy (australopithecus):

above

OpenStudy (australopithecus):

on a piece of paper dont skip out on rules you wont make mistakes if you do, then when you feel comfortable skip them

OpenStudy (australopithecus):

one step at a time

OpenStudy (anonymous):

What do you mean? I don't know how to solve for x.

OpenStudy (australopithecus):

your answer is right

OpenStudy (anonymous):

So would I divide 378.24 by 96.7 and that would be my answer?

OpenStudy (australopithecus):

no, your answer is x

OpenStudy (australopithecus):

x is the volume you are looking for

OpenStudy (anonymous):

So my answer is 96.7?

OpenStudy (australopithecus):

cm^3 is a volume

OpenStudy (australopithecus):

remember units! 378.24g/96.7cm^3 = 3.91 grams/1cm^3

OpenStudy (australopithecus):

do you understand now?

OpenStudy (anonymous):

Yes I do.

OpenStudy (australopithecus):

also remember units cancel out when divided and they increase power when multiplied cm*cm = cm^2 cm/cm = no units cm^2/cm = cm

OpenStudy (australopithecus):

ok I'm gone hope this was insightful take care